OpenStudy (anonymous):

how to identfy and draw the following graph 4x^2+4y^2+8y-3=0

OpenStudy (anonymous):

u'll need some intuition.

OpenStudy (anonymous):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] ellipse

OpenStudy (anonymous):

\[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] hyperbola

OpenStudy (anonymous):

and since the coefficients of the square terms are the same, it is a circle!

OpenStudy (anonymous):

you probably need the center and radius too. do you know how to find them?